A box remains at rest on an incline. The force of
Answer #1

giveb,

mg = 10 N

theta = 40 degree

normal force = mg * cos(40)

normal force = 10 * cos(40)

normal force = 7.66 N

static friction force = mg * sin(40), since the box is in rest

static friction force = 6.428 N

static friction force = coefficient of friction * normal force

6.428 = coefficient of friction * 7.66

coefficient of friction = 0.839